OpenStudy (anonymous):

A wheel contains eight slices, numbered sequentially 1 through 8. The probability of landing on each slice is equal. If the wheel is spun three times, what is the probability that the first spin is 4 or less, the second spin is 6 or less, and the third spin is 8 or less?

OpenStudy (kropot72):

First spin: The probability of 4 is 1/8, the probability of 3, 2 and 1 are also 1/8 for each. These four possible events are mutually exclusive. Therefore we get: \[P(4\ or\ less)=\frac{1}{8}+\frac{1}{8}+\frac{1}{8}+\frac{1}{8}=you\ can\ calculate\]